# Optimal Selection Using Algorithmic Rankings with Side Information

### v2

@@ 5.2 Welfare for candidates:
Next, we will consider impacts that the firm’s hiring strategy has on candidate welfare. For example, a firm using the first-busy strategy will select busy candidates much more frequently than a firm using the first-free strategy. Picking a busy candidate could involve negative impacts: for example, it could be viewed as firms “poaching” candidates from other firms rather than hiring unemployed candidates, which may lead to a higher unemployment rate. However, picking a busy candidate could also be viewed positively: candidates who receive competing offers could potentially use them to increase their compensation.
@@
Increasing the accuracy of the ranking tool increases the probability that the picked candidate is busy with among firms using first-free and may decrease the probability that the picked candidate is busy among firms using first-busy.
Figure 2 illustrates this phenomenon with the RUM with Gumbel noise model. Specifically, both figures include colored, dashed lines showing the probability of a firm picking a busy candidate for every possible window length. Note that for both plots, as the accuracy increases, the colored lines increase: holding window length constant, increasing accuracy increases how frequently busy candidates are picked. For the top figure (first free), as the window width increases, the probability of picking a busy candidate decreases, but for the bottom figure (first-busy), the probability decreases with window width: this is simply because shrinking the window size makes it less likely that each firm will find their preferred free/busy type within the window. The solid black lines in Figure 2 shows how the firm’s strategies change in respond to the ranking tool’s increase in accuracy. Recall from Lemma 2 that increased accuracy always shrinks window size. If a firm is using first-free, then increased accuracy causes the firm to move up and to the right along the colored lines, always increasing how often busy candidates are picked. However, if a firm is using first-busy, then increasing accuracy causes the firm’s strategy to move down and to the right - and ultimately could increase or decrease the probability that a busy candidate is picked, depending on the degree to which the accuracy changes.
Figure 2: Figures showing how changing the accuracy of the tool changes how frequently the picked candidate is busy. The top shows a firm using “first free” strategy: note that increasing accuracy uniformly increases the probability that the picked candidate is busy. The bottom figure shows a firm using first-busy strategy: note that increasing accuracy could increase or decrease the probability that the picked candidate is busy. Parameters: $v_{1}=1,v_{2}=0,p_{1}=0.1,p_{2}=0.4$, each point run with $10^{6}$ simulations. First-free strategy run with $\gamma=10$, first-busy run with $\gamma=1.6$.
@@ 5.3 Strategic use of ranking tool (welfare of ranking company):
Finally, we discuss the impact of increased accuracy on how often firms end up using the ranking tool as it was designed - specifically, on how often firms select the top-ranked candidate. In light of Theorem 3, one might be tempted to assume that increased accuracy leads to higher “congestion at the top,” in the form of firms being more likely to select highly-ranked candidates even if they are busy. While it is true that this can occur, the relationship between accuracy and firm selection can also be more subtle: there are instances in which increasing the accuracy of the ranking tool actually decreases the likelihood of picking the top-ranked candidate.
@@ 6.1 Welfare
For the beyond-the-superstar setting, if the firm is using a first-free or first-busy strategy with a window size that weakly shrinks with increased accuracy, the following properties hold: 1) If the firm is using first-free, then increased accuracy increases the probability that the picked candidate is busy. 2) If the firm is using first free, then increased accuracy may decrease the probability that the picked candidate is busy.
@@ 6.2.1 Restricting choice to either top-free or top-busy candidate
The following permutation distributions are inversion monotone: 1) Mallows, 2) RUM with Gumbel noise, and 3) RUM with noise from any identical, symmetric, noise distribution (e.g. Gaussian).
@@ 6.2.2 Algorithm for picking candidate
1 do
2   Define $r_{i,k}=p_{k}\cdot(1-p_{i})/((1-p_{k})\cdot p_{i})$.
3   Define $\sigma_{i,k,j}$ as permutation with every item ordered correctly, but with $v_{i}$ ranked 1st and $v_{k}$ is ranked $j$th.
4   Define $\tilde{}\sigma_{i,k,j}=\sigma_{i,k,j}$ except with indices $1,j$ flipped.
@@
5 then
6    $G_{ik}=r_{i,k}\cdot P[\sigma]\cdot(v_{k}-v_{i}/\gamma_{i})+P[\tilde{}\sigma]\cdot(v_{i}-v_{k}/\gamma_{k})$
7
8 else
@@
9    $G_{ik}=r_{i,k}\cdot P[\sigma]\cdot(v_{k}/\gamma_{k}-v_{i})+P[\tilde{}\sigma]\cdot(v_{i}/\gamma_{i}-v_{k})$
10
@@
if $\left|G_{j}\right|>\left|G_{1}\right|$
@@ Acknowledgments
This work was partially supported by a MIT METEOR Fellowship, and was partially completed while Kate Donahue was at an internship at Microsoft Research. We wish to thank the following individuals and groups for their very helpful comments: Yeganeh Alimohammadi, Keegan Harris, John Horton, Meena Jagadeesen, Sloan Neitert, Manish Raghavan, Michelle Si, Kiran Tomlinson, and Nicolas Wu, and attendees the following events: Boston College Computer Science Colloquium, Simons Collaboration on the Theory of Algorithmic Fairness Annual Meeting, Themis Science Talk at Amazon, the AI and Science Group at Meta, Harvard Economics and Computer Science Colloquium, INFORMS Annual Meeting, and the AImpactCenter at UIUC for their valuable discussions and insightful suggestions. We also thank the anonymous reviewers for their constructive feedback.
Figure 3: Figure showing examples where increasing accuracy could lead to increased probability of picking the top ranked candidate for first-free strategies (top figure) or decreased for first-busy (bottom). In both cases, phenomeona occurs for very low accuracy (far left part of plot). Both figures show $N=10$ with only lines shown for $i\in[1,4]$ to show detail, with RUM with Gumbel noise. Top has $p_{1}=0.01<p_{2}=0.05$, bottom has $p_{1}=0.9,p_{2}=0.95$. Numerical simulations with $10^{6}$ simulations per point.
Figure 4: Simulation of performance of strategies with $N=4$. Each line gives the expected utility of different strategies, where “best” gives the Bayes-optimal posterior strategy and Alg gives the performance of Algorithm 2 (which reduces to Algorithm 1 in the superstar setting). The shaded regions illustrate where different named strategies happen to also be optimal: for example, the purple region is one where “follow the ranking” is exactly optimal. Unshaded regions are those where the optimal strategy cannot be described as $k$-free or $k$-busy. However, note that the error (gap between yellow and green lines) is often extremely small. The top and middle are both superstar settings, while the bottom is beyond-the-superstar. Note that the top has 0 error always because $v_{2}=0$ (Theorem 1), while the middle has nonzero error. The bottom (beyond-the-superstar) has regions with non-zero error, but largely follows the optimal strategy.
Appendix A Numerical Simulations
@@
For $\gamma=1.5$, as in Figure 6, we have $R=4.3$. By Algorithm 1, this means that the firm should follow “first busy”, and given the normalized values $P[\sigma^{1}]/P[\sigma^{i}]$, the window size should be $k=3$ (because $k=3$ is the maximum index such that $P[\sigma^{1}]/P[\sigma^{i}]<4.3$. Figure 6 numerically simulates the expected utility for every top-k strategy, both first-free and first-busy, showing that first-busy with window size $k=3$ is, in fact, optimal!
@@
Finally, for $\gamma=4$, as in Figure 8, we end up with $R=1.63$, which by Algorithm 1 results in a first-busy strategy with window size $k=1$ - which is equalivant to “follow-the-ranking”. Figure 8 shows that for this setting, the simulation similarly shows that this strategy is optimal.
@@ B.1 Properties of ranking permutations
[From [plackett1975analysis], [luce1959individual]] Consider sets of items with values $v_{i}$, and a Random Utility Model with i.i.d. Gumbel noise $G(\mu,\beta)$ is added to each value. Then, the probability of observing the ordering $\sigma$ is given by:
@@
$\prod_{j=1}^{k-1}\frac{\exp(v_{2}/\beta)}{\exp(v_{1}/\beta)+(N-j)\cdot\exp(v_{2}/\beta)}\cdot\frac{\exp(v_{1}/\beta)}{\exp(v_{1}/\beta)+(N-k)\cdot\exp(v_{2}/\beta)}\cdot\prod_{j=k}^{N}\frac{\exp(v_{2}/\beta)}{\exp(v_{2}/\beta)\cdot(N-j)}$
$\prod_{j=1}^{k-1}\frac{1}{\exp((v_{1}-v_{2})/\beta)+(N-j)}\cdot\frac{\exp((v_{1}-v_{2})/\beta)}{\exp((v_{1}-v_{2})/\beta)+(N-k)}\cdot\frac{1}{\prod_{j=k}^{N}(N-j)}$
$\frac{\exp((v_{1}-v_{2})/\beta)}{\exp((v_{1}-v_{2})/\beta)+N-1}\cdot\prod_{j=2}^{k}\frac{1}{\exp((v_{1}-v_{2})/\beta)+N-j}\cdot\frac{1}{\prod_{j=k}^{N}(N-j)}$
@@
$\frac{\exp((v_{1}-v_{2})/\beta)}{\exp((v_{1}-v_{2})/\beta)+N-1}\cdot\prod_{j=1}^{k-1}\frac{1}{\exp((v_{1}-v_{2})/\beta)+N-j-1}\cdot\frac{1}{\prod_{j=k}^{N}(N-j)}$
@@
$\frac{(N-1)!}{\prod_{j=k}^{N}(N-j)}=\prod_{j=1}^{k-1}(N-j)$
@@
We can observe that the model in Proposition 1 satisfies $P[\sigma^{1}]^{\prime}>P[\sigma^{1}]^{\prime}$, so it must be that $\prod_{k=1}^{i-1}q_{k}^{\prime}<\prod_{k=1}^{i-1}q_{k}$. Because $q_{i}^{\prime}\leq q_{i}<1$, we must have that $\prod_{k=1}^{i}q_{k}^{\prime}<\prod_{k=1}^{i}q_{k}$, which implies that
@@
The most common noise distribution with the Random Utility Model is Gaussian noise, while our theoretical results (Proposition 1) uses Gumbel noise as in the Placket-Luce distribution ([plackett1975analysis, luce1959individual]). Here, we show that empirical results for Gaussian noise closely mimic theoretical results for Gumbel noise. Figure 11 shows PDFS for Gumbel and Gaussian distributions, Figure 12 shows the $P[\sigma^{i}]$ values for the superstar model, and Figure 13 shows the ratio $P[\sigma^{1}]/P[\sigma^{i}]$, both for the Gumbel and Gaussian noise distribution. In all settings, the $\beta$ and standard deviation values are set to be equal.
@@ B.2 General proofs for Section 4
Figure 14: Illustration of optimal strategy for firm, given $N=5$ candidates with $v_{2}=0$ and permutations given by the RUM with Gumbel noise model). The x axis varies the $\gamma$ busy penalty, while the y axis varies accuracy as parameterized by the $1/\beta$ Gumbel noise parameter (higher values increase accuracy). Shades of red indicate regions where first-free selection is the optimal strategy and shades of blue indicate where first-busy selection is (darker shades indicate regions where the firm has larger $j^{*}$ or $k^{*}$ and thus “hunts” further down the list to find a candidate with their preferred free/busy status. The purple region is where the optimal strategy is “follow the ranking”. This figure uses $r=2$, so $\gamma=2$ is the critical point on the x-axis where firms switch from using first-free selection to first-busy selection.
@@ Appendix C Proofs for Section 5
Note: all proofs in this section only rely on accuracy properties as stated in Definition 2: See 2
@@
If necessary, we then then repeat the process with the new minimum index $\ell^{\prime}$ such that $P[\rho^{\ell^{\prime}}]>P[\sigma^{\ell^{\prime}}]$ until we arrive at $P[\rho^{k+1}]^{\prime}=P[\sigma^{k+1}]$.
@@
Note that $p_{1}-p_{2}<0$, so if the term inside the parantheses is positive, then the entire righthand side is negative, and this quantity automatically holds. If the term inside the parantheses is negative, we can rearrange the righthand side:
@@
even though $P[\sigma^{1\leq i\leq k}]<\cdot P[\sigma^{1\leq i\leq k+1}]$. We can use the fact that:
@@
Next, we consider the case that a firm is using first-free with window size $k$. A firm picks the top either either when a) the first candidate is free, or b) all top $k$ candidates are busy. What is probability that this happens?
@@ D.1 Inversion-monotone properties (optimality of picking first free or first busy)
Our main theorem is given by: See 7 We will prove this theorem through a series of smaller lemmas.
Lemma 11.
@@
$v_{2}\cdot(p_{1}\cdot p_{2}\cdot(1-p_{2})\cdot(1-\epsilon)+p_{2}^{2}\cdot(1-p_{1})\cdot\epsilon)>\gamma\cdot v_{1}\cdot p_{2}^{2}\cdot(1-p_{1})\cdot\epsilon$
$v_{2}\cdot(p_{1}\cdot(1-p_{2})\cdot(1-\epsilon)+p_{2}\cdot(1-p_{1})\cdot\epsilon)>\gamma\cdot v_{1}\cdot p_{2}\cdot(1-p_{1})\cdot\epsilon$
@@
$v_{2}\cdot(p_{1}\cdot(1-p_{2})\cdot(1-\epsilon)+p_{2}\cdot(1-p_{1})\cdot\epsilon)>\gamma\cdot v_{1}\cdot p_{2}\cdot(1-p_{1})\cdot\epsilon$
@@
Definition 3.
A probability distribution over permutations $\mathcal{P}$ is called inversion-monotone if it satisfies the following condition: for any permutation $\sigma$ with a pair of indices $i<j$ such that $v_{\sigma_{i}}>v_{\sigma_{j}}$, we construct a corresponding permutation $\tilde{\sigma}$ where $\tilde{\sigma}=\sigma$ except for indices $i,j$, which we have flipped: $v_{\tilde{\sigma}_{i}}=v_{\sigma_{j}},v_{\tilde{\sigma}_{j}}=v_{\sigma_{i}}$. Then, we require:
$P[\sigma]\geq P[\tilde{\sigma}$]
Lemma 12.
Consider any vector of realized status vector $s$, with $s_{i}=s_{j}$, for some $i<j$. Then, if the probability distribution of permutations is inversion-monotone as in Definition 3, $\mathbb{E}[v_{\sigma_{i}}\mid s]\geq\mathbb{E}[v_{\sigma_{j}}\mid s]$, which implies that the optimal solution will always be to pick the first free item or the first busy item.
@@
First, Lemma 13 shows that the Mallows model is inversion-monotone. While helpful, this is not extremely surprising: the Mallows model [mallows1957non] is constructed such that the probability of a permutation $\sigma$ occurring is directly tied to the number of pairs of items that are inverted, so this property follows naturally.
@@
While the Mallows model is frequently used as a model of permutations, it does have drawbacks. Specifically, one desirable property of rankings is that items with very different true values should be less likely to be inverted. For example, given $v_{1}=10,v_{2}=9,v_{3}=0]$, we would expect items $v_{1},v_{2}$ to be more frequently inverted than $v_{2},v_{3}$, even though their ordinal ranks differ by the same amount. This type of property cannot be expressed by the Mallows model, but it can be captured by the Random Utility Model [thurstone1994law]. As described in Section 2, in the Random Utility Model (RUM), while each item has some true value $\{v_{i}\}$, it is assumed that they are ranked by noised versions of these values, such as additive noise ($\hat{}v_{i}\sim\mathcal{D}$). This automatically satisfies the property that items with more similar true values will be more likely to be swapped. Additionally, it seems likely that the Random Utility Model might better capture the performance of rankings produced by humans or algorithmic tools, which might have some sense of the “true value” of each item, but make small, independent errors in their estimation of these values. However, does the Random Utility Model satisfy inversion-monotonicity?
@@
These results tell us that, even in settings beyond the superstar model, for cases where permutations are generated by commonly-used models, the optimal strategy will still be to either pick the first free item or the first busy item.
@@
Consider any permutation $\sigma$ with at least one inversion (an instance where $v_{\sigma_{i}}<v_{\sigma_{j}}$ but $i<j$). Such a permutation occurs when $\hat{}v_{\sigma_{i}}>\hat{}v_{\sigma_{j}}$. By Lemmas 14 and 15, we know that given $v_{\sigma_{i}}<v_{\sigma_{j}}$ and pair of noised values $\hat{}v_{i},\hat{}v_{j}$ generated with with noise either a) from a Gumbel distribution or b) any additive symmetric noise distribution, if $\left|\hat{}v_{\sigma_{i}}-\hat{}v_{\sigma_{j}}\right|=\Delta$, then
@@
First, note that the mean of a Gumbel distribution is given by $\mu+\beta\cdot\eta$, where $\eta$ here is Euler’s constant. That means that we can model $X_{i}=\mu_{i}+\epsilon_{i}$ as $X_{i}\sim G(X_{i}-\beta\cdot\eta)$. Next, we consider the difference distribution induced by $\epsilon_{1}-\epsilon_{2}$: we note that the difference distribution between two Gumbel distributions is given by a Logistic distribution governed by mean $X_{1}-\beta\cdot\eta-X_{2}+\beta\cdot\eta=X_{1}-X_{2}$ and noise parameter $\beta$. Because $X_{1}>X_{2}$, this distribution is right-skewed: thus, for all values $x$, the probability of $X_{1}>X_{2}$ given $\left|X_{1}-X_{2}\right|=x$ is always greater than the probability of $X_{1}<X_{2}$, as desired. ∎
See 15

### v3

@@ 5.2 Welfare for candidates:
Next, we will consider impacts that the firm’s hiring strategy has on candidate welfare. For example, a firm using the first-busy strategy will select busy candidates much more frequently than a firm using the first-free strategy. Picking a busy candidate could involve negative impacts, particularly in terms of equity: for example, it could be viewed as firms “poaching” candidates from other firms rather than hiring unemployed candidates, which may lead to a higher unemployment rate. However, picking a busy candidate could also be viewed positively: candidates who receive competing offers could potentially use them to increase their compensation.
@@
Increasing the accuracy of the ranking tool increases the probability that the picked candidate is busy among firms using first-free and may decrease the probability that the picked candidate is busy among firms using first-busy.
Figure 2 illustrates this phenomenon with the RUM with Gumbel noise model. Specifically, both figures include colored, dashed lines showing the probability of a firm picking a busy candidate for every possible window length. Note that for both plots, as the accuracy increases, the colored lines increase: holding window length constant, increasing accuracy increases how frequently busy candidates are picked. For the left figure (first-free), as the window width increases, the probability of picking a busy candidate decreases, but for the right figure (first-busy), the probability decreases with window width: this is simply because shrinking the window size makes it less likely that each firm will find their preferred free/busy type within the window. The solid black lines in Figure 2 shows how the firm’s strategies change in response to the ranking tool’s increase in accuracy. Recall from Lemma 2 that increased accuracy always shrinks window size. If a firm is using first-free, then increased accuracy causes the firm to move up and to the right along the colored lines, always increasing how often busy candidates are picked. However, if a firm is using first-busy, then increasing accuracy causes the firm’s strategy to move down and to the right - and ultimately could increase or decrease the probability that a busy candidate is picked, depending on the degree to which the accuracy changes.
Figure 2: Figures showing how changing the accuracy of the tool changes how frequently the picked candidate is busy (fixed window sizes dashed, optimal window size solid). The top shows a firm using “first-free” strategy: note that increasing accuracy uniformly increases the probability that the picked candidate is busy. The bottom figure shows a firm using first-busy strategy: note that increasing accuracy could increase or decrease the probability that the picked candidate is busy. Parameters: $v_{1}=1,v_{2}=0,p_{1}=0.1,p_{2}=0.4$, each point run with $10^{6}$ simulations. First-free strategy run with $\gamma=10$, first-busy run with $\gamma=1.6$.
@@ 5.3 Strategic use of ranking tool (welfare of ranking company):
Finally, we discuss the impact of increased accuracy on how often firms end up using the ranking tool as it was designed - specifically, on how often firms select the top-ranked candidate.
In light of Theorem 3, one might be tempted to assume that increased accuracy leads to higher “congestion at the top,” in the form of firms being more likely to select highly-ranked candidates even if they are busy. While it is true that this can occur, the relationship between accuracy and firm selection can also be more subtle: there are instances in which increasing the accuracy of the ranking tool actually decreases the likelihood of picking the top-ranked candidate.
@@ 6.1 Welfare
For the beyond-the-superstar setting, if the firm is using a first-free or first-busy strategy with a window size that weakly shrinks with increased accuracy, the following properties hold: 1) If the firm is using first-free, then increased accuracy increases the probability that the picked candidate is busy. 2) If the firm is using first-busy, then increased accuracy may decrease the probability that the picked candidate is busy.
@@ 6.2.1 Restricting choice to either top-free or top-busy candidate
The following permutation distributions are inversion monotone: 1) Mallows, 2) RUM with Gumbel noise, and 3) RUM with noise from any identical, symmetric noise distribution (e.g. Gaussian).
@@ 6.2.2 Algorithm for picking candidate
do
   Define $r_{i,k}=p_{k}\cdot(1-p_{i})/((1-p_{k})\cdot p_{i})$.
   Define $\sigma_{i,k,j}$ as permutation with every item ordered correctly, but with $v_{i}$ ranked 1st and $v_{k}$ is ranked $j$th.
   Define $\tilde{}\sigma_{i,k,j}=\sigma_{i,k,j}$ except with indices $1,j$ flipped.
@@
  then
    $G_{ik}=r_{i,k}\cdot P[\sigma]\cdot(v_{k}-v_{i}/\gamma_{i})+P[\tilde{}\sigma]\cdot(v_{i}-v_{k}/\gamma_{k})$
 else
@@
1    $G_{ik}=P[\sigma]\cdot(v_{k}/\gamma_{k}-v_{i})+r_{i,k}\cdot P[\tilde{}\sigma]\cdot(v_{i}/\gamma_{i}-v_{k})$
2
@@
if $\left|\mathcal{G}_{j}\right|>\left|\mathcal{G}_{1}\right|$
@@ Acknowledgments
This work was partially supported by a MIT METEOR Fellowship, and was partially completed while Kate Donahue was at an internship at Microsoft Research. We wish to thank the following individuals and groups for their very helpful comments: Yeganeh Alimohammadi, Keegan Harris, John Horton, Meena Jagadeesen, Sloan Neitert, Manish Raghavan, Michelle Si, Kiran Tomlinson, and Nicolas Wu, and attendees of the following events: Boston College Computer Science Colloquium, Simons Collaboration on the Theory of Algorithmic Fairness Annual Meeting, Themis Science Talk at Amazon, the AI and Science Group at Meta, Harvard Economics and Computer Science Colloquium, INFORMS Annual Meeting, and the AImpactCenter at UIUC for their valuable discussions and insightful suggestions. We also thank the anonymous reviewers for their constructive feedback.
Appendix A Numerical extensions
@@
For $\gamma=1.5$, as in Figure 6, we have $R=4.3$. By Algorithm 1, this means that the firm should follow “first-busy”, and given the normalized values $P[\sigma^{1}]/P[\sigma^{i}]$, the window size should be $k=4$ (because $k=4$ is the maximum index such that $P[\sigma^{1}]/P[\sigma^{i}]<4.3$. Figure 6 numerically simulates the expected utility for every top-k strategy, both first-free and first-busy, showing that first-busy with window size $k=3$ is, in fact, optimal!
@@
Finally, for $\gamma=4$, as in Figure 8, we end up with $R=1.63$, which by Algorithm 1 results in a first-busy strategy with window size $k=1$ - which is equivalent to “follow-the-ranking”. Figure 8 shows that for this setting, the simulation similarly shows that this strategy is optimal.
@@ B.1 Properties of ranking permutations
[From [21], [16]] Consider sets of items with values $v_{i}$, and a Random Utility Model with i.i.d. Gumbel noise $G(\mu,\beta)$ is added to each value. Then, the probability of observing the ordering $\sigma$ is given by:
@@
$\prod_{j=1}^{k-1}\frac{\exp(v_{2}/\beta)}{\exp(v_{1}/\beta)+(N-j)\cdot\exp(v_{2}/\beta)}\cdot\frac{\exp(v_{1}/\beta)}{\exp(v_{1}/\beta)+(N-k)\cdot\exp(v_{2}/\beta)}\cdot\prod_{j=k}^{N-1}\frac{\exp(v_{2}/\beta)}{\exp(v_{2}/\beta)\cdot(N-j)}$
$\prod_{j=1}^{k-1}\frac{1}{\exp((v_{1}-v_{2})/\beta)+(N-j)}\cdot\frac{\exp((v_{1}-v_{2})/\beta)}{\exp((v_{1}-v_{2})/\beta)+(N-k)}\cdot\frac{1}{\prod_{j=k}^{N-1}(N-j)}$
$\frac{\exp((v_{1}-v_{2})/\beta)}{\exp((v_{1}-v_{2})/\beta)+N-1}\cdot\prod_{j=2}^{k}\frac{1}{\exp((v_{1}-v_{2})/\beta)+N-j}\cdot\frac{1}{\prod_{j=k}^{N-1}(N-j)}$
@@
$\frac{\exp((v_{1}-v_{2})/\beta)}{\exp((v_{1}-v_{2})/\beta)+N-1}\cdot\prod_{j=1}^{k-1}\frac{1}{\exp((v_{1}-v_{2})/\beta)+N-j-1}\cdot\frac{1}{\prod_{j=k}^{N-1}(N-j)}$
@@
$\frac{(N-1)!}{\prod_{j=k}^{N-1}(N-j)}=\prod_{j=1}^{k-1}(N-j)$
@@
We can observe that the model in Proposition 1 satisfies $P[\sigma^{1}]^{\prime}>P[\sigma^{1}]$, so it must be that $\prod_{k=1}^{i-1}q_{k}^{\prime}<\prod_{k=1}^{i-1}q_{k}$. Because $q_{i}^{\prime}\leq q_{i}<1$, we must have that $\prod_{k=1}^{i}q_{k}^{\prime}<\prod_{k=1}^{i}q_{k}$, which implies that
@@
The most common noise distribution with the Random Utility Model is Gaussian noise, while our theoretical results (Proposition 1) uses Gumbel noise as in the Placket-Luce distribution ([21, 16]). Here, we show that empirical results for Gaussian noise closely mimic theoretical results for Gumbel noise. Figure 11 shows PDFS for Gumbel and Gaussian distributions, Figure 12 shows the $P[\sigma^{i}]$ values for the superstar model, and Figure 13 shows the ratio $P[\sigma^{1}]/P[\sigma^{i}]$, both for the Gumbel and Gaussian noise distribution. In all settings, the $\beta$ and standard deviation values are set to be equal.
@@ Appendix C Proofs for Section 5
Note: all proofs in this section only rely on accuracy properties as stated in Definition 3: See 3
@@
If necessary, we then repeat the process with the new minimum index $\ell^{\prime}$ such that $P[\rho^{\ell^{\prime}}]>P[\sigma^{\ell^{\prime}}]$ until we arrive at $P[\rho^{k+1}]^{\prime}=P[\sigma^{k+1}]$.
@@
Note that $p_{1}-p_{2}<0$, so if the term inside the parentheses is positive, then the entire righthand side is negative, and this quantity automatically holds. If the term inside the parentheses is negative, we can rearrange the righthand side:
@@
even though $P[\sigma^{1\leq i\leq k}]<P[\sigma^{1\leq i\leq k+1}]$. We can use the fact that:
@@
Next, we consider the case that a firm is using first-free with window size $k$. A firm picks the top candidate either when a) the first candidate is free, or b) all top $k$ candidates are busy. What is probability that this happens?
@@ D.1 Inversion-monotone properties (optimality of picking first free or first busy)
Our main theorem is given by: See 7 We will prove this theorem through a series of smaller lemmas. See 4
@@
$v_{2}\cdot(p_{1}\cdot p_{2}\cdot(1-p_{2})\cdot(1-\epsilon)+p_{2}^{2}\cdot(1-p_{1})\cdot\epsilon)>1/\gamma\cdot v_{1}\cdot p_{2}^{2}\cdot(1-p_{1})\cdot\epsilon$
$v_{2}\cdot(p_{1}\cdot(1-p_{2})\cdot(1-\epsilon)+p_{2}\cdot(1-p_{1})\cdot\epsilon)>1/\gamma\cdot v_{1}\cdot p_{2}\cdot(1-p_{1})\cdot\epsilon$
@@
$v_{2}\cdot(p_{1}\cdot(1-p_{2})\cdot(1-\epsilon)+p_{2}\cdot(1-p_{1})\cdot\epsilon)>1/\gamma\cdot v_{1}\cdot p_{2}\cdot(1-p_{1})\cdot\epsilon$
@@
See 2
See 5
@@
First, Lemma 11 shows that the Mallows model is inversion-monotone. While helpful, this is not extremely surprising: the Mallows model [17] is constructed such that the probability of a permutation $\sigma$ occurring is directly tied to the number of pairs of items that are inverted, so this property follows naturally.
@@
While the Mallows model is frequently used as a model of permutations, it does have drawbacks. Specifically, one desirable property of rankings is that items with very different true values should be less likely to be inverted. For example, given $v_{1}=10,v_{2}=9,v_{3}=0]$, we would expect items $v_{1},v_{2}$ to be more frequently inverted than $v_{2},v_{3}$, even though their ordinal ranks differ by the same amount. This type of property cannot be expressed by the Mallows model, but it can be captured by the Random Utility Model [23]. As described in Section 2, in the Random Utility Model (RUM), while each item has some true value $\{v_{i}\}$, it is assumed that they are ranked by noised versions of these values, such as additive noise ($\hat{}v_{i}\sim\mathcal{D}$). This automatically satisfies the property that items with more similar true values will be more likely to be swapped. Additionally, it seems likely that the Random Utility Model might better capture the performance of rankings produced by humans or algorithmic tools, which might have some sense of the “true value” of each item, but make small, independent errors in their estimation of these values. However, does the Random Utility Model satisfy inversion-monotonicity?
@@
These results tell us that, even in settings beyond the superstar model, for cases where permutations are generated by commonly-used models, the optimal strategy will still be to either pick the first free item or the first busy candidate.
@@
Consider any permutation $\sigma$ with at least one inversion (an instance where $v_{\sigma_{i}}<v_{\sigma_{j}}$ but $i<j$). Such a permutation occurs when $\hat{}v_{\sigma_{i}}>\hat{}v_{\sigma_{j}}$. By Lemmas 12 and 13, we know that given $v_{\sigma_{i}}<v_{\sigma_{j}}$ and pair of noised values $\hat{}v_{i},\hat{}v_{j}$ generated with with noise either a) from a Gumbel distribution or b) any additive symmetric noise distribution, if $\left|\hat{}v_{\sigma_{i}}-\hat{}v_{\sigma_{j}}\right|=\Delta$, then
@@
First, note that the mean of a Gumbel distribution is given by $\mu+\beta\cdot\eta$, where $\eta$ here is Euler’s constant. That means that we can model $X_{i}=\mu_{i}+\epsilon_{i}$ as $X_{i}\sim G(X_{i}-\beta\cdot\eta)$. Next, we consider the difference distribution induced by $\epsilon_{1}-\epsilon_{2}$: we note that the difference distribution between two Gumbel distributions is given by a Logistic distribution governed by mean $X_{1}-\beta\cdot\eta-X_{2}+\beta\cdot\eta=X_{1}-X_{2}$ and noise parameter $\beta$. Because $X_{1}>X_{2}$, this distribution is shifted towards the right: thus, for all values $x$, the probability of $X_{1}>X_{2}$ given $\left|X_{1}-X_{2}\right|=x$ is always greater than the probability of $X_{1}<X_{2}$, as desired. ∎
See 13